OpenStudy (anonymous):

show that dy/dx= x+2/ 2squareroot 1+x u=x v=(1+x)-1/2 u`=1 v`=-1/2(1+X)^-3/2

OpenStudy (asadkarim7):

u ( v') + v ( u')

OpenStudy (anonymous):

its u ( v') - v ( u') but I cant get the answer it give 3x+2/ 2squareroot(1+x)

OpenStudy (asadkarim7):

oh wait

OpenStudy (asadkarim7):

its v (u') - u (v')/ v^2

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

but if you use u=x v=(1+x)-1/2 u`=1 v`=-1/2(1+X)^-3/2 I cant get it to give i3x+2/ 2squareroot(1+x)

OpenStudy (asadkarim7):

let me try this out

OpenStudy (asadkarim7):

can u post the original equation ?

OpenStudy (anonymous):

the original equation is y=x square root 1+X then you have to show that dy/dy= 3x+2/2 suare root 1+X

OpenStudy (asadkarim7):

oh k lets try this 1

OpenStudy (asadkarim7):

i got the answer right

OpenStudy (anonymous):

could u please show your solution becuase I think my working out is wrong

OpenStudy (asadkarim7):

1st of all u = x v = sqrt 1+x u' = 1 v' = 1/2sqrt 1+x formula u (v') + v( u') x ( 1/2 sqrt 1+x) + sqrt 1+x ( 1) x/ 2 sqrt 1+x + sqrt 1+x take lcm x + 2 (sqrt( 1+x )) (sqrt 1+x )/ 2 sqrt 1+x x + 2 ( 1+x) / 2 sqrt 1+x x+ 2 +2x / 2sqrt 1+x 3x + 2 / 2 sqrt 1+x

OpenStudy (anonymous):

Thankyou :)

OpenStudy (asadkarim7):

N p :) where r u from ?

OpenStudy (anonymous):

Rochdale But I go Oldham Sixth Form wbu?

OpenStudy (asadkarim7):

Pakistan :)

OpenStudy (anonymous):

oh lol Im in England :)

OpenStudy (asadkarim7):

yeah i know :P

OpenStudy (anonymous):

how come you usedu (v') + v( u') when the formlua is u (v') - v( u') over \[2^{(v)}\]

OpenStudy (anonymous):

I mean \[\over v)^{2}\]

OpenStudy (asadkarim7):

the formula u ( v') + v ( u') is used when in the equation u and v are multiplied v ( u') - u ( v') / V^2 is used when u and v are divided in the equation.

OpenStudy (asadkarim7):

in this equation x is multiplied with sqrt x+1

OpenStudy (asadkarim7):

so we used 1st formula !

OpenStudy (anonymous):

ohright I get it now :) thanks lol where about do u live in pakistan?

OpenStudy (asadkarim7):

Karachi :) and u muslim ?

OpenStudy (anonymous):

yeah :)

OpenStudy (asadkarim7):

Masha' Allah :D

OpenStudy (anonymous):

Thankyou :)

OpenStudy (anonymous):

how did you get? x/ 2 sqrt 1+x + sqrt 1+x and what do you mean by lcm?

OpenStudy (asadkarim7):

you need to multiply the other term with the denominator to get the single denominator.